# IFAdapter: Instance Feature Control for Grounded Text-to-Image Generation

### v1

@@ 1 Introduction
The advent of diffusion models has revolutionized the field of Text-to-Image (T2I) synthesis  (Ho et al., 2020; Podell et al., 2023; Baldridge et al., 2024; Betker et al., 2023; Rombach et al., 2022; Yang et al., 2023a). Despite their exceptional performance in generating high-quality images of single instances, these models remain limited in generating images containing multiple instances. There are two key challenges underscore this limitation: 1) The ambiguity of sequential natural language descriptions of spatial information impedes the accurate transmission of user intent to the model, resulting in the inability to precisely position instances. 2) Relying solely on a global prompt to describe the features of multiple instances fails to correctly associate these features with their corresponding instances (Feng et al., ).
@@
Recent advancements in the Layout-to-Image (L2I) task (Li et al., 2023; Wang et al., 2024c; Zhou et al., 2024b; Kim et al., 2023; Bar-Tal et al., 2023) have enabled diffusion models to achieve precise instance-level position control by incorporating bounding boxes as spatial signals. However, in terms of instance feature generation, most state-of-the-art (SOTA) L2I methods can accurately depict only coarse features of an instance (e.g., color attribution), while struggling to generate more complex, fine-grained features, as shown in Figure 1(a). This shortcoming limits the models’ applicability in scenarios such as graphic design and fashion design, where a high level of detail in local regions is essential. To this end, we propose a more challenging task, termed Instance Feature Generation (IFG) task, which aims to simultaneously address both the positional accuracy and feature generation accuracy of instances. We found that existing T2I methods do not perform satisfactorily on the IFG task. Upon analysis, we attribute this phenomenon to two factors: 1) Training Data: Most L2I methods rely solely on category labels as descriptions for instances during training. This approach causes samples with detailed descriptions to become out-of-distribution during inference. 2) Insufficient Feature Information: Existing designs mostly use a single contextualized token to guide the feature generation of each instance. While this token effectively captures the coarse semantics of the instance (Chen et al., 2024), it is limited in generating high-frequency appearance features.
@@
In this work, we propose the Instance Feature Adapter (IFAdapter) to address the aforementioned limitations. First, to address issues related to the training data, we utilize existing SOTA Vision-Language Models (VLMs) for annotation, generating a dataset with detailed instance-level descriptions. Subsequently, we implement two meticulously designed components to address the challenges of instance positioning and feature representation. 1) Appearance Tokens: To address the loss of detailed feature information in instances, the IFAdapter introduces novel learnable appearance queries. These queries extract instance-specific feature information from descriptions, forming appearance tokens that work alongside $EoT$ tokens, thereby enabling more precise control over the generation of instance features; 2) Instance Semantic Map: In contrast to sequence-to-2D grounding conditions (Li et al., 2023; Wang et al., 2024c), IFAdapter constructs a 2D semantic map to correlate instance features with designated spatial locations. This map-like condition provides enhanced spatial guidance, reinforcing the spatial prior and preventing the leakage of instance features. In regions where multiple instances overlap, a gated semantic fusion mechanism is employed to resolve feature confusion. The IFAdapter integrates the semantic map only within a subset of cross-attention layers (Vaswani, 2017) in the diffusion model. This loose coupling allows the IFAdapter to function as a plug-and-play component, enabling its instance-level control capabilities to be transferred across various community models without requiring retraining, as illustrated in 1(b).
@@ 4.6 Ablation Study
Figure 5: Ablation results.
@@
75.8 66.5 58.8 48.3 25.0 22.0
Table 3: Results of ablation study.

### v2

@@ 1 Introduction
The advent of diffusion models has revolutionized the field of Text-to-Image (T2I) synthesis  (Ho et al., 2020; Podell et al., 2023; Baldridge et al., 2024; Betker et al., 2023; Rombach et al., 2022; Yang et al., 2023a). Despite their exceptional performance in generating high-quality images of single objects, these models remain limited in composing multiple objects into an exquisite image. There are two key challenges underscore this limitation: 1) The inability of natural language in conveying precise spatial information impedes expression of user intent to the model, resulting in poor image composition in the generated images. 2) Relying solely on a given text prompt describing the attributes of multiple objects, existing models often fails to bind the detailed features to the correct object instances (Feng et al., ).
@@
Recent advancements in the Layout-to-Image (L2I) task (Li et al., 2023; Wang et al., 2024c; Zhou et al., 2024b; Kim et al., 2023; Bar-Tal et al., 2023) have partially mitigated such limitation and achieved precise instance-level position control by incorporating bounding boxes as spatial signals. However, in terms of instance feature generation, most state-of-the-art (SOTA) L2I methods can only accurately depict coarse features of an instance (e.g., color attribution), while struggling to generate more complex, fine-grained features, as shown in Figure 1(a). This shortcoming limits the models’ applicability in scenarios such as graphic design and art design, where local high-grade details are essential. To simultaneously track the improvement of layout accuracy and feature generation accuracy, we introduce a more challenging task, termed Instance Feature Generation (IFG) task. We found that existing T2I methods do not perform satisfactorily on the IFG task. Upon experiment and analysis, we observe that existing T2I methods do not perform satisfactorily on the IFG task and ascribe this phenomenon to two restrictions: 1) Insufficient detailed descriptions: Most L2I methods rely solely on category labels as descriptions for instances during training. This approach causes samples with detailed descriptions to become out-of-distribution during inference. 2) Insufficient feature information: Existing designs mostly use a single contextualized token to guide the feature generation of each instance. Although this token effectively captures the coarse semantics of the instance (Chen et al., 2024), it is limited in generating high-frequency appearance features.
@@
In this work, we propose the Instance Feature Adapter (IFAdapter) to address the aforementioned restrictions. First, to address issues related to the training data, we utilize existing SOTA Vision-Language Models (VLMs) for annotation, generating a dataset with detailed instance-level descriptions. Subsequently, we implement two meticulously designed components to address the challenges of instance positioning and feature representation. 1) Appearance Tokens: To address the loss of detailed feature information in instances, the IFAdapter introduces novel learnable appearance queries. These queries extract instance-specific feature information from descriptions, forming appearance tokens that work alongside $EoT$ tokens, thereby enabling more precise control over the generation of instance features; 2) Instance Semantic Map: In contrast to sequence-to-2D grounding conditions (Li et al., 2023; Wang et al., 2024c), IFAdapter constructs a 2D semantic map to correlate instance features with designated spatial locations. This map-like condition provides enhanced spatial guidance, reinforcing the spatial prior and preventing the leakage of instance features. In regions where multiple instances overlap, a gated semantic fusion mechanism is employed to resolve feature confusion. The IFAdapter integrates the semantic map only within a subset of cross-attention layers (Vaswani, 2017) in the diffusion model. This loose coupling allows the IFAdapter to function as a plug-and-play component, enabling its instance-level control capabilities to be transferred across various community models without requiring retraining, as illustrated in 1(b).
@@ 4.6 Ablation Study
Figure 5: Qualitative results of variants of IFAdapter.
@@
79.7 68.6 61.0 49.0 25.1 22.0
Table 3: Quantitative results of variants of IFAdapter.